[clas_members] [New Logentry] RC Daily Summary

daria at jlab.org daria at jlab.org
Sat Sep 10 12:50:02 EDT 2022

Logentry Text:

Hello RG-C people, 

Things are generally going well. 

Yesterday, Sergey and Nathan implemented the read-out of the helicity-latched triggers and the calculation of the helicity asymmetry (thank you!!) -- there's a new GUI for this now and a new set of strip-charts, being saved to the log-book for every run. See details and a discussion of the values here: https://logbooks.jlab.org/entry/4039705

We ran with the positive target polarisation until about 6.15pm last night, then switched target polarisation with Maxwell guiding the process over the phone. Since the microwave & NMR controls are temperamental (need to be attended to during the Wednesday maintenance day), shifts to the frequency are done manually now, from the controller in the electronics room in the counting house. Initial polarisation change is quick, then the increase to a "peak" value is pretty slow. It took about 2.5h before we were happy that we found the microwave sweet-spot: 140.48 GHz. 

We started taking data (run 16923) when polarisation reached -20%. It gradually increased in amplitude during this run, reaching -29.6% at 38M events and -31% at the end of the run (105M). 

During the night we continued to take production data with HW plate IN and negative polarisation on ND3. 

A couple of glitches and a pre-dawn call to the RICH expert (thank you Valery!): 
* RICH communication issue needing a soft reboot, then later in the night some tiles crashing and needing Valery's intervention to bring them back up. 
* A DAQ crash at the end of a run.

Day running well -- except a logbook issue (there are two Day Shift Summaries now).

This morning the NMR is reading -31%. Current plan -- unless there are objections or dragons emerge from analysis of the data that Mohammed is cooking (thank you!) -- is to continue running as long as polarisation doesn't drop, ideally to the end of the weekend, then anneal on Monday. This evening we will flip the HW plate. 

We will NOT have an RC meeting today unless new info emerges and someone calls for one! Same regarding Sunday :-)

Have a good weekend!

Your so-far-quite-chilled RC.

******** Stats since the start of ND3 with ELMO ********
Half-wave plate OUT, target polarisation POSITIVE:

16902: 100.4 M (pinned)
16903: 102 M
16904: 100.1 M (pinned)
16905: 121.8 M
16907: 42.6 M (DAQ crash at end)
16908: 37.4 M (presumed junk! missing part of FTOF)

TOTAL triggers from "good" runs: 467 M

Half-wave plate IN, target polarisation POSITIVE:

16910: 91.1 M
16911: 101.6 M
16912: 71.8 M (beamline devices disconnected at the end)
16913: 86.4 M
16918: 41.2 (not clear whether pinned)
16922: 106.0 M
TOTAL triggers from "good" runs: 498 M (slightly lower target pol than the set with OUT half-wave plate).

Half-wave plate IN, target polarisation NEGATIVE:

16923: 105.3 M (target pol increasing during this run).
16924: 100.0 M (pinned)
16925: 44.0 M (Brief RICH glitch, DAQ crash at the end)
16927: 100.0 M
16928: running... currently 93 M... next run will be pinned.

TOTAL triggers from "good" runs so far: 442 M


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4039941
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220910/d8dcc37e/attachment.html>

More information about the clas_members mailing list